What is the rate constant and rate constant units for a 3.5 order reaction?

Given a rate of 5.2x10^-3 M/s; and the rate orders 1, 1/2, and 2 (given all three reactant concentrations are 0.10M).

I got 16.44 M^-2.5 s^-1 but I'm not sure if my units are correct.
Answer #1

Answer :

Rate Law or Rate Equation is the equation that links the reaction rate with the concentration of reactants or pressures of reactants.

According to Law of Mass Action,Rate Of Reaction is directly proportional to the concentration of reactants.

In this example consider reaction as,

A + B + C = Products

Rate Law becomes,Rate = K [A]^x[ B]^y[C]^z,

Rate Orders,x=1,y=0.5,z=2................ Given

Rate= 5.2×10^-3 M/s (given),

Order Of Reaction = 3.5( Given),

Concentration of each reactant= 0.10 M (given)

So to find rate constant put all Value in rate equation,Rate= K [0.10]^1 ×[0.10]^0.5 ×[0.10]^2 ,

5.2 × 10^-3 = K [3.16 × 10^-4] ,

K =5.2 ×10^-3/3.16 ×10^-4

K = 16.443 M^-2.5 s^-1 (Rate Constant)

Now to calculate unit, put units in rate equation,

Ms^-1 = K [M]^1+0.5+2

= K [M]^3.5  

K = Ms^-1 / M^3.5

K = M^-2.5 s^-1(Unit Of Rate Constant)